Circular No. 5803

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1993h) Carolyn S. Shoemaker, Eugene M. Shoemaker, and David H. Levy re-

port their discovery of another comet. The following positions are available:

1993 UT  α_{2000}  δ_{2000}  m_{1} Observer
May 23.20885 13^{h}24.^{m}8  −34^{o}00′  16.5 Shoemaker
24.20572 13 24.0  −34 01  _{”}
25.18776 13 23 14.^{s}45 −34 00 21′′.3 16.7 Scotti
25.20008 13 23 13.82 −34 00 21.8  _{”}
25.21155 13 23 13.27 −34 00 22.0 16.4  _{”}
25.39850 13 23 04.61 −34 00 31.0  McNaught

C. S. Shoemaker, E. M. Shoemaker, and D. H. Levy (Palomar). 0.46-m Schmidt telescope films. Comet moderately condensed; tail 15′′ long in p.a. 20^{o}.

J. V. Scotti (Kitt Peak). 0.9-m Spacewatch telescope. Coma diameter 16′′;

narrow tail extends 2′.14 in p.a. 5^{o}; m_{2} = 20.8 on May 25.20. R. H. McNaught (Siding Spring). Uppsala Southern Schmidt telescope.

V1974 CYGNI G. S. Stringfellow and S. Bowyer, Center for Extreme Ultraviolet As-

trophysics, University of California at Berkeley, report: “V1974 Cyg (N Cyg 1992) was detected by the Extreme Ultraviolet Explorer (EUVE) during the course of the all-sky survey. V1974 Cyg was scanned by the EUVE between 1992 Nov. 26 and Dec. 9, by which time it had faded to V ∼ 10. The count rate and associated uncertainty in the Lexan/Boron filter was 0.110 ± 0.011 counts/s. This bandpass is down from peak sensitivity by about an order of magnitude at 6.0 nm and a factor of 50 at 3.0 nm. If V1974 Cyg lies at a distance > 0.5 kpc, then the interstellar N(H) is expected to be of order 1.0 × 10^{21} cm^{−2}, which should provide complete self-absorption of the EUV flux. Comparison of the above count rates with those of Sco X-1, also detected by EUVE, suggests that the EUV detections result from soft x-rays. This implies that V1974 Cyg had entered into a coronal phase by late November 1992. The soft x-ray flux detected by EUVE from V1974 Cyg rivals that of Sco X-1, making it one of the brightest soft x-ray sources in the sky at that time.”

1993 May 25 Daniel W. E. Green